Calculate the number of moles in 19.0 g C6H12O6
Answer #1

la U LA ST Molecular weight of CH1206. 6CC) + 12CH) + 6(0) = 6612) + 12(1)+6016) z 72+12 + 6 = 180 Molecular weight of COM120